OpenStudy (anonymous):

y = 7/3x^3 rewrite, differentiate, simplify (without using the quotient rule)

OpenStudy (anonymous):

y=7/3 * x^-3 dy/dx = (7/3)*-3*x^-4 =-7/(x^4)